# Multi-frequency analysis of the ALMA and VLA high resolution continuum observations of the substructured disc around CI Tau

### v1

@@ 5.3 Result comparison for different dust mixtures
Jäger ($\,{}^{\circ}$) 0.17 0.36 0.03 0.44 ($\,{}^{\circ}$)
Jäger ($\,{}^{\circ}$) 0.17 0.36 0.03 0.44 ($\,{}^{\circ}$)
Jäger ($\,{}^{\circ}$) 0.17 0.36 0.03 0.44 ($\,{}^{\circ}$)
Jäger ($\,{}^{\circ}$) 0.17 0.36 0.03 0.44 ($\,{}^{\circ}$)
@@
 [^11] Slightly different bulk densities for (astro)silicates ($3.50\,{\rm g}\,{\rm cm}^{-3}$) and water ($1.00\,{\rm g}\,{\rm cm}^{-3}$) are used in dsharp_opac (Birnstiel et al., 2018) to generate the optical properties for the “Ricci (compact)” mixture. Zubko et al. (1996) do not report any bulk density for their amorphous carbonaceous grains; a density of $2.50\,{\rm g}\,{\rm cm}^{-3}$ is adopted in dsharp_opac (Birnstiel et al., 2018) following Ricci et al. (2010b). Note, however, that Woitke et al. (2016) and Dominik et al. (2021) use $1.80\,{\rm g}\,{\rm cm}^{-3}$ for the BE-sample. This difference has negligible effects on the optical properties computed with dsharp_opac. Bulk dust mixture densities are: $1.70\,{\rm g}\,{\rm cm}^{-3}$ for “Ricci (compact)”, $1.68\,{\rm g}\,{\rm cm}^{-3}$ for “DSHARP (default)”, $2.11\,{\rm g}\,{\rm cm}^{-3}$ for “DSHARP (icefree)”, $1.40\,{\rm g}\,{\rm cm}^{-3}$ for “DSHARP (highwater)”, $2.12\,{\rm g}\,{\rm cm}^{-3}$ for “Zubko (ACH2)” and “Zubko (BE)”, $1.83\,{\rm g}\,{\rm cm}^{-3}$ for “Jäger”, $2.01\,{\rm g}\,{\rm cm}^{-3}$ for “Draine” opacities.
@@
First of all, for our fiducial dust composition, Figure 6 shows that $\chi^{2}\approx 1$ with a small uncertainty ($\lesssim 1$ for $R\lesssim 82\,{\rm au}$), thus confirming (cf. Appendix F) that “Ricci (compact)” grains can fit remarkably well CI Tau’s surface brightness and spectral index radial profiles at most radii. We compared our fiducial results with those obtained using different dust optical properties, starting from the default mixture adopted by the DSHARP collaboration (Andrews et al., 2018; Birnstiel et al., 2018), labelled “DSHARP (default)”. In this case, while the brightness profiles are fitted reasonably well through all the disc ($\chi^{2}_{\rm SB}\approx 2$), $\alpha_{\rm B7-B6}$ is systematically underestimated, leading to a larger $\chi^{2}_{\rm SI}\approx 3$ for the spectral index radial profiles. This difference from our fiducial results does not depend on the different fraction of water ice adopted in the two mixtures, that only marginally affects dust optical properties (as already noticed by Birnstiel et al. 2018). Indeed, increasing the dust water ice content to 60% by volume (as in the “Ricci (compact)” composition) or removing it altogether (compositions labelled “DSHARP (highwater)” and “DSHARP (icefree)”, respectively), no clear difference from the “DSHARP (default)” results can be seen in Figure 6, neither for the surface brightness, nor for the spectral index radial profiles.
Instead, the different organic materials (hence dielectric constants) adopted by Ricci et al. (2010b) and Birnstiel et al. (2018) impact our results the most. Depending on their synthesis temperature (e.g., Jäger et al. 1998), interstellar carbonaceous material can come in a mixture of different hybridisation states, such as sp2 (where sheets of carbon atoms are formed, as in graphite), sp3 (where carbon atoms are arranged in tetrahedral geometry, as in diamonds), or even mixed states (where curved structures, such as fullerene, are abundant). The relative fraction of these hybridisation states and the amount of hydrogenation heavily affect the dielectric constants of carbon grains (e.g., Zubko et al. 1996), with substantial effects on their opacities. To test the sensitivity of our results to these properties, following Birnstiel et al. (2018), we generate optical constants for dust mixtures with the same volume fractions as the “DSHARP (default)” ones, but replacing the refractory organics with different carbonaceous materials [^12] See Appendix B of Birnstiel et al. (2018), reference notebook here.. First, we tested the dielectric constants of amorphous carbon grains produced by arc discharge in hydrogen atmosphere (ACH2-sample) or burning benzene (BE-sample) proposed by Zubko et al. (1996), labelled “Zubko (ACH2)” and “Zubko (BE)”, respectively. In both cases, Figure 6 shows that our models are in excellent agreement with the data, with $\chi^{2}\approx 1$ and a small spread ($\lesssim 1$ for $R\lesssim 82\,{\rm au}$) both for the surface brightness and spectral index radial profiles. Noticeably, as reported in subsection 4.1 and footnote 11, also our fiducial “Ricci (compact)” opacities were generated using the dielectric constants of Zubko et al. (1996) for ACH2 amorphous carbon, further strengthening the evidence that amorphous carbon grains can fit our data remarkably well. The different fractions of mixing materials adopted in the “Ricci (compact)” and “Zubko (ACH2)” compositions, instead, suggests that (i) small fractions of troilite do not substantially affect our results, and (ii) our observations cannot discriminate between mixtures of the same materials in different fractions (within $\lesssim 15\%$), among those tested in Figure 6 (see footnote 11).
Instead, when the carbonaceous materials in the mixtures are made mostly by graphite, our models provide poorer fits to the data. This can be seen in Figure 6 when the refractory organics (Henning & Stognienko, 1996) in the “DSHARP (default)” mixture are replaced by carbonaceous materials synthesised by pyrolising cellulose at progressively higher temperatures (Jäger et al. 1998, labelled “Jäger” followed by the pyrolysis temperature). Indeed, while organics produced by low temperature pyrolysis are made essentially by amorphous carbon and fit our data reasonably well ($\chi^{2}\approx 2$ to 3), at temperature higher than $\geq 800\,^{\circ}$ such grains become dominated by large graphitic areas and fit our data much worse ($\chi^{2}\gg 5$). Such low quality fits were also obtained adopting the graphite optical constants of (Draine 2003, labelled “Draine” followed by the carbon grain size and the relative direction of the electric field and the graphite plane our dielectric constants were computed for), confirming our hypothesis that mixtures including organics made by amorphous carbon can fit our data the best.
Such different results for different compositions can be explained by the Mie interference location and magnitude in their absorption opacity spectral index. In the case of “Ricci (compact)”, “Zubko”, and “DSHARP (default)” compositions, such a resonance leads to a peak-like opacity spectral index profile, with $\beta^{\rm abs}>3$ at $a_{\rm max}\approx\lambda/2\pi$ for $q\leq 2.5$. Instead, for “Jäger (¿800 C)” and “Draine” compositions, the opacity spectral index is sigmoid-shaped and $\leq 2.5$, regardless of $q$. The number of resonances present varies from none up to a few, but they are low-amplitude and just make the $\beta^{\rm abs}$ profile look more jagged. Since dust emission is optically thin in VLA Ka Band (regardless of the compositions adopted in our analysis), only the former set of opacities can explain a $\beta_{\rm B3-Ka}=\alpha_{\rm B3-Ka}-2>2.5$ for $R\gtrsim 30\,{\rm au}$ (see Figure 3). New (sub-)cm wavelength observations deep enough to detect dust emission with high SNR across the entire disc could provide stronger indications that $\alpha_{\rm B3-Ka}>5$ in the outer disc. This will be useful to further constrain dust composition because different peak magnitudes in the absorption spectral index can be seen between BE- and AC2H-sample amorphous carbon species (e.g., $\max(\beta^{\rm abs})\approx 4.3$ and 4.7 for “Zubko (ACH2)” and “Ricci (compact)”, while $\max(\beta^{\rm abs})\approx 3.3$ for “Zubko (BE)”), but also for different water ice volume fractions (e.g., $\max(\beta^{\rm abs})\approx 3.9$ for “DSHARP (icefree)” opacities, while $\max(\beta^{\rm abs})\approx 3.1$ for the “DSHARP (highwater)” ones).
For the “Ricci (compact)”, “Zubko (BE)”, and “DSHARP (default)” compositions we also performed high angular resolution fits as detailed in subsection 4.3. Their marginalised posterior distributions are discussed in Appendix G and indicate that (i) the surface density and maximum grain size retrieved for the “Ricci (compact)” and “Zubko (BE)” mixtures are consistent within a factor of three, and (ii) the “DSHARP (default)” mixture gives less reliable results.
To summarise, CI Tau’s curse of a steepening spectral index between ALMA Band 3 and VLA Ka Band, that makes it difficult to detect and resolve dust emission at wavelengths longer than a few millimetres, turned out to be its blessing, because only very few dust mixtures, those with the sharpest Mie resonance transition, such as “Ricci (compact)” and “Zubko”, can explain this feature, providing unexpectedly strong constraints on the bulk dust composition in this disc.
5.4 Result comparison for different porosities
We tested the impact of dust porosity on our results fitting the low resolution surface brightness radial profiles with “Ricci” and “DSHARP” dust compositions, but progressively increasing their void fraction from 10%, 30%, 50%, and 70% up to 90% by volume [^13] We focused on the mixtures that best reproduce CI Tau’s observations for compact grains. Since “Ricci” and both “Zubko” mixtures give similar results, hereafter we only discuss the former.. We used the dsharp_opac package (Birnstiel et al., 2018) to determine the dielectric constants of each mixture with the Bruggeman rule [^14] Here we followed Zhang et al. (2023), while Birnstiel et al. (2018) preferred the Maxwell-Garnett rule (Maxwell Garnett, 1904) for porous grains, treating the physical mixture as inclusion in a background “void matrix”. Ricci et al. (2010b) also used the Bruggeman rule for their default optical properties (i.e., “Ricci” mixture with 30% porosity). These two methods lead to significant differences in the optical constants, and a generally worse agreement with the data when the Maxwell-Garnett rule is adopted. Furthermore, following Kataoka et al. (2014) and the argument of Ueda et al. (2024), we treated porous particles as aggregates of compact monomers, mixing first the refractive materials, then adding void as a separate component, as explained here. and their optical properties using Mie theory for spherical grains. As for the comparison of different dust mixtures (subsection 5.3), we kept the temperature fixed while fitting for the other parameters. Overall, for both the “Ricci” and “DSHARP” compositions we noticed an increase of the best fit density and grain size with the porosity fraction, especially for filling factors $>50\%$ and $>30\%$ for the “Ricci” and “DSHARP” grain mixtures, respectively. This trend can be motivated by the reduced absorption opacity and the progressive shift of the Mie resonance to sizes much larger than the reference wavelength for fluffier particles.
Figure 7: Reduced $\chi^{2}$ between our model posteriors and the observed surface brightness (upper panel) and spectral index (bottom panel) radial profiles for different porosity fractions, colour-coded by the outermost disc radius the $\chi^{2}$ was averaged over. Dots and diamonds are used for the “Ricci” and “DSHARP” mixtures, respectively. The shaded grey regions indicate where the inferred dust surface density is high enough for the disc to be in the marginally gravitational unstable regime under the assumption of a standard dust-to-gas ratio $Z=10^{-2}$. The “Ricci” mixture provides better fits than the “DSHARP” one for every porosity level and suggests that dust in CI Tau can be up to 50% porous. The “DSHARP” mixture, instead, can explain the data only for compact aggregates.
To estimate the quality of these fits, we measured the reduced $\chi^{2}$ between our models and the surface brightness and spectral index radial profiles following the procedure introduced in subsection 4.4. The median and $1\sigma$ uncertainty of the reduced $\chi^{2}$ posterior distributions as a function of the adopted porosity fraction, colour-coded by the outermost disc radius the $\chi^{2}$ was computed for, are displayed in Figure 7. The top and bottom panels summarise our model agreement with the surface brightness and spectral index radial profiles, respectively. Our fits with the “Ricci” mixture (red, grey, and blue dots) can reproduce remarkably well both the brightness and spectral index profiles for compact grains and particles with porosity fractions up to 50%. In these cases, even though the estimated $\chi^{2}$ slightly increases with the void fraction, it remains below two. For every chosen porosity, the “DSHARP” mixture (pink, grey, and green diamonds) fits our data worse than the “Ricci” one, with $\chi^{2}>3$ already for void fractions as small as 10%. For both “Ricci” and “DSHARP” mixtures, porosities $>50\%$ lead to very poor quality fits. In addition for highly porous particles (with void fractions of 70% and 90% or larger, for the “DSHARP” and “Ricci” composition, respectively) our inferred dust density is so high that, for a standard dust-to-gas ratio $Z=10^{-2}$, CI Tau should be marginally gravitationally unstable (Toomre parameter $Q<1$, grey shaded area in Figure 7). Since we observe no clear evidence of instability in the dust and gas emission, we favour lower porosity fractions.
The differences between “Ricci” and “DSHARP” results can be attributed once again to shape of the $\beta^{\rm abs}$ profile. Increasing the porosity fraction in the dust mixture leads to progressively shallower Mie resonances in the absorption opacity spectral index (for a fixed $q$), that almost completely disappears for void fractions $>50\%$. Therefore, the “DSHARP” composition gives worse results than the “Ricci” one because, for any given porosity, its optical properties lead to much shallower $\beta^{\rm abs}$ profiles, so shallow that even reducing $q$ (as in Figure 17) does not provide enough flexibility to compensate for the effect of porosity on the Mie resonance. A similar justification applies to the case of “Ricci” opacities with $>70\%$ porosity.
To summarise, dust porosity fractions up to $\approx 50\%$ are compatible CI Tau’s observations when considering “Ricci” opacities. The “DSHARP” mixture, instead, provides good quality fits almost only for compact particles. This result suggests that, by robustly assessing the particle porosity (e.g., using full polarisation observations, see subsection 6.2), we can tell “Ricci” and “DSHARP” opacity (or ultimately amorphous carbonaceous or refractory organic) models apart.
6 Discussion
6.1 Interpretation of the best fit radial profiles
Maximum grain size
A puzzling result from subsection 5.1 is our smooth and almost constant best fit grain size radial profile.

### v2

@@ Appendix I Absorption spectral index
[…]
@@
 [^24] Slightly different bulk densities for (astro)silicates ($3.50\,{\rm g}\,{\rm cm}^{-3}$) and water ($1.00\,{\rm g}\,{\rm cm}^{-3}$) are used in dsharp_opac (Birnstiel et al. 2018) to generate the optical properties for the ‘Ricci (compact)’ mixture. Zubko et al. (1996) do not report any bulk density for their amorphous carbonaceous grains; a density of $2.50\,{\rm g}\,{\rm cm}^{-3}$ is adopted in dsharp_opac (Birnstiel et al. 2018) following Ricci et al. (2010b). We note, however, that Woitke et al. (2016) and Dominik et al. (2021) use $1.80\,{\rm g}\,{\rm cm}^{-3}$ for the BE-sample. This difference has negligible effects on the optical properties computed with dsharp_opac. Bulk dust mixture densities are: $1.70\,{\rm g}\,{\rm cm}^{-3}$ for ‘Ricci (compact)’, $1.68\,{\rm g}\,{\rm cm}^{-3}$ for ‘DSHARP (default)’, $2.11\,{\rm g}\,{\rm cm}^{-3}$ for ‘DSHARP (icefree)’, $1.40\,{\rm g}\,{\rm cm}^{-3}$ for ‘DSHARP (highwater)’, $2.12\,{\rm g}\,{\rm cm}^{-3}$ for ‘Zubko (ACH2)’ and ‘Zubko (BE)’, $1.83\,{\rm g}\,{\rm cm}^{-3}$ for ‘Jäger’, $2.01\,{\rm g}\,{\rm cm}^{-3}$ for ‘Draine’ opacities.
@@
Table 6: Comparison between pebble masses measured in the optically thin approximation and from multi-frequency continuum observations for the best studied sources in the literature and CI Tau.
Name Flux density (1.3 mm) Distance Pebble mass (OT) Pebble mass (MF) Outer radius Composition Comments References
(mJy) (pc) ($M_{\rm Earth}$) ($M_{\rm Earth}$) (au)
(1) (2) (3) (4) (5) (6) (7) (8) (9)
HL Tau 744.1 147 441.89 332.95 100 Pollack $q=3.5$ 1, 2, 3
TW Hya 575.4 60.1 57.19 250 to 330 70 DSHARP (default) Large grains 4, 5, 4
210 to 310 Large grains, $q=3.5$
IM Lup 253 153.8 153.62 $1211.92^{+662.56}_{-469.45}$ 150 DSHARP (default) Large grains, $q=2.5$ 7, 6, 8
GM Aur 264.1 155.0 129.13 $246.38^{+113.20}_{-43.28}$ 175 DSHARP (default) 9, 6, 8
AS 209 288 120.4 107.16 $249.71^{+239.72}_{-69.92}$ 125 DSHARP (default) Large grains, $q=2.5$ 7, 6, 8
HD 163296 715 103.7 197.46 $276.35^{+449.48}_{-59.93}$ 120 DSHARP (default) Large grains, $q=2.5$ 7, 6, 8
265 DIANA 7, 5, 10
MWC 480 268 156.2 179.74 $396.21^{+589.31}_{-96.55}$ 100 DSHARP (default) Large grains, $q=2.5$ 11, 5, 8
LkCa 15 418.1 154.8 292.05 253.12 120 DSHARP (default) $q=3$ 12, 6, 12
9.99 Ricci (compact)
CI Tau 147.7 160.3 107.50 $31.13^{+8.77}_{-6.00}$ 200 Ricci (compact) 13, 6, 13
$28.25^{+8.34}_{-5.58}$ Zubko (BE)
$250.40^{+51.46}_{-37.73}$ DSHARP (default)
 [^25] Column 1: Source name. Column 2: ALMA Band 6 flux density. Column 3: Source distance. Column 4 and 5: Pebble mass in the optically thin approximation (Equation 11, OT) and the surface density profile from multi-frequency observations (MF). Column 6: Outermost disc radius the surface density was integrated over. Column 7 and 8: Adopted composition and fitting assumptions. DIANA opacities (Min et al. 2016; Woitke et al. 2016, labelled ‘DIANA’), are made of 60% amorphous silicates (Dorschner et al. 1995) and 15% amorphous carbonaceous materials (Zubko et al. 1996, BE-sample), with 25% porosity by volume. The mixture was generated using the Bruggeman rule, with bulk density $\rho_{\rm s}=2.08\,{\rm g}\,{\rm cm}^{-3}$. Carrasco-González et al. (2019) generated a set of opacities adopting the optical constant of Pollack et al. (1994, labelled ‘Pollack’). Their mixture is made of 26% silicates, 31% organics, and 43% water ice by volume, with bulk density $\rho_{\rm s}=2.05\,{\rm g}\,{\rm cm}^{-3}$.
[…]
Flux density, distance, and pebble mass from: (1) ALMA Partnership et al. (2015); (2) Galli et al. (2018); (3) Carrasco-González et al. (2019); (4) Mac´ıas et al. (2021); (5) Gaia Collaboration (2020); (6) Gaia Collaboration et al. (2023); (7) Andrews et al. (2018); (8) Sierra et al. (2021); (9) Huang et al. (2020a); (10) Guidi et al. (2022); (11) Liu (2019); (12) Sierra et al. (2025); (13) this paper.